# CustomVideo: Customizing Text-to-Video Generation with Multiple Subjects

### v2

@@ 3.3 Model Training and Inference
During inference, CustomVideo only requires a specific text prompt with corresponding learned word token integrated to generate a required video. To be note that ground truth masks of subjects are not required during inference. Although the model is fine-tuned with single-frame videos extended from subject images, our proposed CustomVideo can generate videos with high diversity while not losing the the fidelity of subjects and motion smoothness (see Figure 1 and Figure 3).
4 Experiment
4.1 Experimental Setup
Dataset.
We collect a dataset CustomStudio with 63 individual subjects and construct 68 meaningful pairs for multi-subject driven T2V generation. These subjects are adapted from DreamBooth [45], CustomDiffusion [40], and Mix-of-Show [52], in which they cover a wide range of 13 categories. Each pair of subjects has 10 different text prompts for evaluation, which are designed with different contexts, actions, and so on. More details of dataset can be found in Sec. A.
Implementation Details.
We train CustomVideo for 500 steps with AdamW [53] optimizer (batch size 2, learning rate 4e-5, and weight decay 1e-2). For class-specific prior preservation, we generate 200 class images with Stable Diffusion v2.1 [13] for each subject, in which the generating prompts are obtained from Claude-3-Opus [54]. Note that the corresponding class images are also concatenated during training phase. The negative value in the mask is set as -1e-8. The weight parameters and in Eq. (4) are set as 0.2 and 1.0, respectively. During inference, we perform 50 steps denoising with DPM-Solver [8] sampler and classifier-free guidance [55]. The resolution of generated 24-frame videos are with 8 fps. Interestingly, we find that the trained weights and word tokens learned from low resolution T2V model can be directly loaded for high resolution zeroscope T2V model [56] to generate personalized videos with resolution, in which no additional training computation cost is required. Our CustomVideo is implemented based on Diffusers [57]. The training phase takes about 8 minutes for a subject pair on 1 RTX 3090 GPU. Meanwhile, it takes about 1 minute and 2 minutes to generate a low and high resolution video on 1 RTX 3090 GPU, respectively.
Comparison Methods.
Except for most relevant VideoDreamer [12], we also consider adapting previous SOTA image-based multi-subject driven methods to a video scenario for comparison, including DreamBooth [45] and CustomDiffusion [40]. Also, we compare our method with single-subject driven video personalization method DreamVideo [11]. Moreover, we consider an alternative way for multi-subject T2V customization, that is, the combination of multi-concept T2I and I2V diffusion models. For this case, we utilize the latest multi-concept driven T2I method DisenDiff [14] for customization and I2V method SVD-XT [15] for animating the generated image.
Evaluation Metrics.
4.2 Main Results
Qualitative Results.
We present the qualitative comparison results in Figure 3. From these results, we observe that our CustomVideo approach effectively ensures the co-occurrence of multiple subjects and successfully disentangles different subjects. However, both DreamBooth and CustomDiffusion fail to capture the structural color information of the car provided in our experiments. The dominant black color from the car window obscures the entire car, resulting in low fidelity. Moreover, the generated frames from VideoDreamer lack consistency, as some frames depict two cars while others only show one car. Additionally, VideoDreamer fails to capture accurate color information, such as the color of the car door. The alternative approach, DisenDiff + SVD-XT can not neither retain the identities of the given subjects nor generate videos with rich motions. Without specific design for tackling multiple subjects, DreamVideo can not distinguish the similar pair ‘cat’ and ‘dog’. In contrast, our CustomVideo method excels in handling such challenging scenes and foreground subject scenarios, effectively capturing the intricate structural details of the provided car for video generation. Similarly, when considering the case of ‘cat’ and ‘dog’, our approach also demonstrates superior capability in generating high-quality videos. More results can be found in Figure 9 and Figure 10.
| Method | Reference | CLIP-T | CLIP-I | DINO-I | T. Cons. |
| DreamBooth [45] | CVPR’23 | 0.6451 | 0.6079 | 0.3109 | 0.7084 |
| CustomDiffusion [40] | CVPR’23 | 0.6524 | 0.6206 | 0.3164 | 0.7342 |
| VideoDreamer [12] | arXiv’23 | 0.6638 | 0.6297 | 0.3479 | 0.7267 |
| DisenDiff [14] + SVD [15] | CVPR’24 | 0.6592 | 0.6179 | 0.3318 | 0.7095 |
| DreamVideo [11] | CVPR’24 | 0.6646 | 0.6128 | 0.3228 | 0.7498 |
| CustomVideo (ours) | / | 0.7075 | 0.6863 | 0.3983 | 0.7960 |
Quantitative Results.
We conduct quantitative experiments on our collected dataset. To ensure a thorough analysis, we generate videos using 10 individual prompts and 4 random seeds for each pair of subjects, resulting in a total of 2,720 generated videos for each method. We then evaluate the quality of the generated videos using four metrics, and the results are presented in Table 1. The table clearly demonstrates that our proposed method is capable of generating videos that are better aligned with the given subjects, outperforming the most recent video diffusion model, DreamVideo, by 11.99% and 23.39% in terms of CLIP Image Alignment and DINO Image Alignment, respectively. Meanwhile, the videos generated from our approach achieve better alignment with the desired text prompts, exceeding DreamVideo by 6.46% CLIP Textual Alignment. These improvements can be attributed to our specially designed co-occurrence and attention control mechanisms, which effectively disentangle and preserve the fidelity of the subjects. Furthermore, our CustomVideo generates videos with significantly higher temporal consistency compared to SOTA methods, as indicated in Table 1. For instance, CustomVideo surpasses DreamVideo by 6.16% in terms of Temporal Consistency.
Human Preference Study.
To further validate our method, we conduct human evaluations on our CustomVideo with comparison to 5 SOTA methods. In this study, we collect 1500 answers from 25 independent human raters with the following questions: 1) which one is aligned to the text prompt best? 2) which one is aligned to the subject images best? 3) which one has the best overall quality? The results are shown in Figure 4. Our CustomVideo is found to be the most preferred option based on human evaluations in all three dimensions.
4.3 Ablation Studies
We conduct ablation analysis of our method from 3 aspects: 1) the effect of each component in our method; 2) the behavior of attention control; 3) the effect of hyper-parameters in attention control.
Component Analysis.
We conduct a thorough analysis of each component in our CustomVideo, presenting both qualitative and quantitative results in Figure 5 and Table 3. One crucial observation is the significance of ensuring co-occurrence during the fine-tuning process. We notice a significant drop in performance when multiple subjects are not concatenated into a single image (line ‘w/o concat’ in Table 3). The absence of subject concatenation leads to the domination of one single subject, which is deemed unacceptable (line ‘w/o concat’ in Figure 5). Moreover, we conduct a study on fine-tuning T2V model with both single and concatenated subjects to examine if the single subject could aid in learning corresponding characteristics. Surprisingly, the results reveal that adding the single subjects to the training process proved detrimental, resulting in inconsistent generation during inference (line ‘both single and concat’ in Figure 5).
Our attention control mechanism plays a crucial role in preserving the identities of subjects. The positive attention guidance in particular improves the CLIP Image Alignment metric by 11.45%, as demonstrated in Table 3. The negative attention guidance also proved beneficial in promoting better image alignment. Moreover, the attention control mechanism significantly enhances the temporal consistency of the generated videos, a crucial factor in video generation. The videos generated without positive and negative attention guidance clearly showcase the positive impact of both components (lines ‘w/o pos. attn.’ and ‘w/o neg. attn.’ in Figure 5). The positive guidance specifically preserves the unique characteristics of a subject, such as distinguishing between a cat and a dog, while the negative guidance weakens the influence of other subjects on a specific subject. By incorporating both positive and negative guidance, our CustomVideo excels in generating videos with high subject fidelity and remarkable temporal consistency. Furthermore, we observe that removing the background from the given subject images significantly improves the generation of high-quality videos (line ‘w/o remove bg’ in Figure 5). By eliminating the background, the T2V model can focus solely on learning the characteristics of the given subjects.
Comparison of Cross-attention Maps.
We investigate the cross-attention maps during training, and study the behavior of our attention control mechanism. The visualization results are shown in Figure 6. Without attention control, it can be observed that the learnable word token ‘<new1>’ for cat is aligned to the region of dog, which is unacceptable. In contrast, equipped with attention control mechanism, the learnable word token ‘<new1>’ and ‘<new2>’ are better aligned to the correct areas of cat and dog, respectively. Such results strongly demonstrate the effectiveness of our approach.
Effect of Hyper-parameters in Attention Control.
We also conduct investigations into the effects of two important parameters: the weight of attention loss () and the negative value () used in the guidance mask (). The quantitative results of these investigations are presented in Table 3a and Table 3b. We observe that the best performance is achieved when the weight is set to 0.2. This value results in optimal alignment between the generated videos and the given subjects, indicating the importance of appropriately balancing the attention guidance during training. For the negative value in the guidance mask, our findings reveal that even a slight negative value is sufficient to enhance the quality of T2V generation. This implies that incorporating negative attention guidance can effectively suppress the influence of other subjects on a specific subject, improving the generation quality.
| Method | CLIP-T | CLIP-I | DINO-I | T. Cons. |
| w/o remove bg | 0.6682 | 0.6752 | 0.3636 | 0.7609 |
| w/o concat | 0.6115 | 0.6611 | 0.3386 | 0.7034 |
| both single and concat | 0.6248 | 0.6096 | 0.3054 | 0.7604 |
| w/o pos. attn. | 0.6851 | 0.6158 | 0.3215 | 0.7213 |
| w/o neg. attn. | 0.6961 | 0.6716 | 0.3613 | 0.7886 |
| CustomVideo (ours) | 0.7075 | 0.6863 | 0.3983 | 0.7960 |
| CLIP-T | CLIP-I | DINO-I | T. Cons. | |
| 1.0 | 0.6909 | 0.6496 | 0.3773 | 0.7812 |
| 0.2 | 0.7075 | 0.6863 | 0.3983 | 0.7960 |
| 0.01 | 0.6924 | 0.6710 | 0.3715 | 0.7886 |
| CLIP-T | CLIP-I | DINO-I | T. Cons. | |
| -1e-5 | 0.6786 | 0.6392 | 0.3719 | 0.7604 |
| -1e-8 | 0.7075 | 0.6863 | 0.3983 | 0.7960 |
| -1e-11 | 0.6926 | 0.6737 | 0.3613 | 0.7812 |
4.4 Limitations
In Figure 7, we present some failure cases encountered during our experiments. Since our method relies on the ability of the base model, it would fail if the base model could not generate, such as small faces in a global view. Currently, Our approach can not tackle too many subjects, in which the spatial positions of different objects in the video conflict with each other. To address such case, we can control the positions of different subjects via conditioning on the pre-defined spatial bounding boxes during inference sampling.
5 Conclusions
Appendix A CustomStudio Dataset
Dataset Info.
We collect a dataset CustomStudio for multi-subject driven video generation. This dataset is composed of 63 individual objects and 68 meaningful pairs. These objects cover a wide range of 13 diverse categories, including anime character, decorate item, furniture, instrument, person, pet, plant, plush, scene, thing, toy, transport, and wearable item. For each subject, 4-21 images are provided as the reference images. The images are adopted from DreamBooth [45], CustomDiffusion [40] and Mix-of-Show [52]. An overview of the samples in CustomStudio dataset is shown in Figure 8. A meaningful pair is composed of different subjects may appear in a video together in practice. For example, ‘cat’ and ‘dog’ can be a meaningful pair, while ‘book’ and ‘barn’ can not arise in the most cases. We set up 25 types of subject pairs, which are constructed based on the corresponding categories, as shown in Table 4.
| Two-subject Pair | [decorate item, plant], [person, transport], [pet, pet], [furniture, thing], [toy, toy] |
| [decorate item, furniture], [instrument, person], [person, person], [plush, plush] | |
| [anime character, anime character], [toy, transport], [person, wearable item] | |
| [person, scene], [scene, transport], [instrument, anime character] | |
| Three-subject Pair | [anime character, anime character, scene], [anime character, instrument, scene] |
| [toy, toy, toy], [plush, plush, plush], [person, instrument, scene] | |
| [person, person, scene], [pet, wearable item, scene], [toy, transport, scene] | |
@@ Appendix A CustomStudio Dataset
Prompt Templates.
The prompt templates used for generating videos during inference are shown in Table 5. For example, we want to generate a customized video with a ‘cat’ and a ‘dog’, the prompt for inference can be ‘a <new1> cat and a <new2> dog walking in the Times Square’.
| Prompt Template for 2-subject Pair |
| a <c1> and a <c2> sitting on an antique table |
| a <c1> and a <c2> sitting on beach with a view of seashore |
| a <c2> and a <c1> side by side on a mountaintop, overlooking a sunrise |
| a <c1> and a <c2> on a surfboard together in the middle of a clear blue ocean |
| a <c1> playing with <c2> |
| a <c1> playing with a robot toy <c2> |
| a <c2> playing with a robot toy <c1> |
| a plush toy replica of a <c1> and a <c2> sitting beside it |
| a <c1> and a <c2> walking in the Times Square |
| a <c1> and a <c2> walking on the Great Wall |
| Prompt Template for 3-subject Pair |
| a <c1> wearing a <c2>, with a <c3> in the background |
| a <c1> wearing a <c2> and giving a speech, with a <c3> in the background |
| a wide shot of a <c1> wearing a <c2> with boston city in background, with a <c3> in the background |
| a long shot of a <c1> walking their golden retriever, wearing a <c2>, with a <c3> in the background |
| a <c1> sitting on the sidewalk wearing a <c2>, with a <c3> in the background |
| a <c1> standing at a graffiti wall, showcasing a new <c2>, with a <c3> in the background |
| close shot of a <c1> walking on a ramp wearing a <c2>, with a <c3> in the background |
| a <c1> stepping out of a taxi in a <c2>, with a <c3> in the background |
| close shot of a <c1> walking in rain wearing a <c2>, with a <c3> in the background |
| long shot of a <c1> sitting on the edge of a roof wearing <c2>, with a <c3> in the background |
Appendix B Additional Experimental Results
Layers of Cross-attention Maps.
We investigates the effect of how we extract the cross-attention maps from the cross-attention layers. There are four levels of cross-attention layers in the U-Net, in which the sizes cross-attention maps are different. Taking the resolution of our generated low resolution video as an example, the sizes of cross-attention maps are (), (), (), and (). Here, we study the effect of different levels of cross-attention maps, the results are shown in Table 6. We observe a significant performance drop with too large () or small () cross-attention maps. Large cross-attention maps may, on one hand, lead to the loss of subtle subject features, while on the other hand, small cross-attention maps can result in inaccurate optimization between the maps and the downsampled ground truth masks. Regarding the cross-attention maps with middle size from and , works better. Thus, in our experiments, we extract cross-attention maps from for attention control.
| CLIP-T | CLIP-I | DINO-I | T. Cons. | ||||
| ✓ | ✗ | ✗ | ✗ | 0.6111 | 0.6335 | 0.3964 | 0.7315 |
| ✗ | ✓ | ✗ | ✗ | 0.6080 | 0.6641 | 0.3055 | 0.7242 |
| ✗ | ✗ | ✓ | ✗ | 0.7075 | 0.6863 | 0.3983 | 0.7960 |
| ✗ | ✗ | ✗ | ✓ | 0.6134 | 0.6420 | 0.3215 | 0.7793 |
| ✓ | ✓ | ✗ | ✗ | 0.6234 | 0.6375 | 0.3497 | 0.7443 |
| ✗ | ✗ | ✓ | ✓ | 0.7047 | 0.6007 | 0.3642 | 0.7380 |
| ✗ | ✓ | ✓ | ✗ | 0.6730 | 0.6364 | 0.3977 | 0.7433 |
| ✓ | ✗ | ✗ | ✓ | 0.6508 | 0.6360 | 0.3055 | 0.7469 |
| ✓ | ✓ | ✓ | ✓ | 0.6432 | 0.6738 | 0.3594 | 0.6975 |
@@ Appendix B Additional Experimental Results
| A majestic Savannah cat with a tall, slender build, surveying its surroundings from atop a bookshelf |
@@
Additional Qualitative Results.
We show additional qualitative results here, including comparison of our CustomVideo with previous SOTA methods in Figure 9 and generated video of CustomVideo in Figure 10. It is clear that our method can generate videos with higher fidelity and better preserve the identities of the reference subjects.
Appendix C Additional Experimental Details
All of the training process is under fp16 mixed precision with accelerate package [62]. We use data augmentations during training, including randomly horizontal flip, randomly crop and resize, together with corresponding prompt change (‘very small’ or ‘close up’ appends before the prompt).
Class Prompts.
For class-specific prior preservation, we generate 200 class images with Stable Diffusion v2.1 [13] for each subject, in which the generating prompts are obtained from Claude-3-Opus [54]. We use Stable Diffusion v2.1 [13] to generate 200 class images for every subject in class-specific prior preservation. The text prompts for generating are obtained from Claude-3-Opus [54]. We show 20 examples in Table 7.
Implementation Details.
We adopt the follow setting for in our experiments, including the previous SOTA methods for comparison and our CustomVideo:
CustomVideo: The weights of query and value in all of the cross-attention layers (including the temporal and spatial layers) of U-Net are fine-tuned. The sampling step for high resolution video is 30.
Human Interface for User Study.
The human annotators are all researchers with experience in artificial intelligence. The interface for human evaluation is shown in Figure 11.
Appendix D Broader Impact
Generating realistic videos with human imagination and personal belongings is crucial in practice. Our work makes this poorly explored problem come true by multi-subject driven T2V generation. We believe our work will benefit a wide range of AIGC applications, such as film production and virtual reality. Although our model significantly outperforms other competitive methods, it still struggle to generate promising results for some cases, such as small faces and lots of subjects shown in Figure 7. Please be careful to use our model in the situations where failures will cause severe consequences.

### v3

@@ III-C Model Training and Inference
During inference, CustomVideo only requires a specific text prompt with corresponding learned word token integrated to generate a required video. To be note that ground truth masks of subjects are not required during inference. Thanks to the thorough optimization of the model and learnable text tokens during training, even challenging scenarios with overlapping subjects—such as the ”car and barn” example—can be effectively managed (see Fig. 1). Although the model is fine-tuned with single-frame videos extended from subject images, our proposed CustomVideo can generate videos with high diversity while not losing the the fidelity of subjects and motion smoothness (see Fig. 1).
IV Experiment
| Two-subject Pair | [decorate item, plant], [toy, toy], [pet, pet], [furniture, thing] |
| [decorate item, furniture], [instrument, person], [person, person] | |
| [anime character, anime character], [person, wearable item] | |
| [plush, plush], [person, transport], [instrument, anime character] | |
| [person, scene], [scene, transport], [toy, transport] | |
| Three-subject Pair | [anime character, anime character, scene], [toy, toy, toy] |
| [anime character, instrument, scene], [plush, plush, plush] | |
| [person, instrument, scene], [person, person, scene] | |
| [pet, wearable item, scene], [toy, transport, scene] | |
@@ IV Experiment
| Prompt Template for Two-subject Pair |
| a [c1] and a [c2] sitting on an antique table |
| a [c1] and a [c2] sitting on beach with a view of seashore |
| a [c2] and a [c1] side by side on a mountaintop, overlooking a sunrise |
| a [c1] and a [c2] on a surfboard together in the middle of a clear blue ocean |
| a [c1] playing with [c2] |
| a [c1] playing with a robot toy [c2] |
| a [c2] playing with a robot toy [c1] |
| a plush toy replica of a [c1] and a [c2] sitting beside it |
| a [c1] and a [c2] walking in the Times Square |
| a [c1] and a [c2] walking on the Great Wall |
| Prompt Template for Three-subject Pair |
| a [c1] wearing a [c2], with a [c3] in the background |
| a [c1] wearing a [c2] and giving a speech, with a [c3] in the background |
| a wide shot of a [c1] wearing a [c2] with boston city in background, with a [c3] in the background |
| a long shot of a [c1] walking their golden retriever, wearing a [c2], with a [c3] in the background |
| a [c1] sitting on the sidewalk wearing a [c2], with a [c3] in the background |
| a [c1] standing at a graffiti wall, showcasing a new [c2], with a [c3] in the background |
| close shot of a [c1] walking on a ramp wearing a [c2], with a [c3] in the background |
| a [c1] stepping out of a taxi in a [c2], with a [c3] in the background |
| close shot of a [c1] walking in rain wearing a [c2], with a [c3] in the background |
| long shot of a [c1] sitting on the edge of a roof wearing [c2], with a [c3] in the background |
@@
| A majestic Savannah cat with a tall, slender build, surveying its surroundings |
@@
IV-A Experimental Setup
Dataset
We collect a dataset CustomStudio for multi-subject driven video generation. This dataset is composed of 63 individual objects and 68 meaningful pairs. These objects cover a wide range of 13 diverse categories, including anime character, decorate item, furniture, instrument, person, pet, plant, plush, scene, thing, toy, transport, and wearable item. For each subject, 4-21 images are provided as the reference images. The images are adopted from DreamBooth [49], CustomDiffusion [43] and Mix-of-Show [57]. An overview of the samples in CustomStudio dataset is shown in Fig.3. A meaningful pair is composed of different subjects may appear in a video together in practice. For example, ‘cat’ and ‘dog’ can be a meaningful pair, while ‘book’ and ‘barn’ can not arise in the most cases. We set up 25 types of subject pairs, which are constructed based on the corresponding categories, as shown in Table I. Each pair of subjects has 10 different text prompts for evaluation, which are designed with different contexts, actions, and so on, as shown in Table II. The text prompts for generating are obtained from Claude-3-Opus [56]. We show 20 examples in Table III.
Implementation Details
We train CustomVideo for 500 steps with AdamW [58] optimizer (batch size 2, learning rate 4e-5, and weight decay 1e-2). For class-specific prior preservation, we generate 200 class images with Stable Diffusion v2.1 [13] for each subject, in which the generating prompts are obtained from Claude-3-Opus [56]. Note that the corresponding class images are also concatenated during training phase. The negative value in the mask is set as -1e-8. The weight parameters and in Eq. (5) are set as 0.2 and 1.0, respectively. During inference, we perform 50 steps denoising with DPM-Solver [8] sampler and classifier-free guidance [59]. The resolution of generated 24-frame videos are with 8 fps. Interestingly, we find that the trained weights and word tokens learned from low resolution T2V model can be directly loaded for high resolution zeroscope T2V model [60] to generate personalized videos with resolution, in which no additional training computation cost is required. The weights of key and value in all of the cross-attention layers (including the temporal and spatial layers) of main blocks within U-Net are fine-tuned. To be note that an individual temporal transformer apart from main blocks within U-Net is responsible for modeling the motion in video generation. This temporal component remains fixed throughout our training process. And we only focus on fine-tuning a small amount of key and value parameters. This targeted optimization allows the model to learn the identities of the provided multiple subjects while leaving the motion generation capability of the base model entirely intact. The sampling step for high resolution video is 30. Our CustomVideo is implemented based on PyTorch [61] and Diffusers [62]. All of the training process is under fp16 mixed precision with accelerate package [63]. We use data augmentations during training, including randomly horizontal flip, randomly crop and resize, together with corresponding prompt change (‘very small’ or ‘close up’ appends before the prompt). The training phase takes about 8 minutes for a subject pair on 1 RTX 3090 GPU. Meanwhile, it takes about 1 minute and 2 minutes to generate a low and high resolution video on 1 RTX 3090 GPU, respectively.
Prompt Templates
The prompt templates used for generating videos during inference are shown in Table II. Taking “a [c1] and a [c2] walking in the Times Square” as an example, “[c1]” and “[c2]” are two positions for integrating the paired class names and learnable text tokens. If we want to generate a video with “cat” and “dog”, we can integrate these names into the above prompt template with the corresponding learned text tokens “new1” and “new2”. To this end, the prompt for inference comes to “a new1 cat and a new2 dog walking in the Times Square” or “a new1 dog and a new2 cat walking in the Times Square”. Both two prompts works for customized generation.
Comparison Methods
Except for most relevant VideoDreamer [12], we also consider adapting previous SOTA image-based multi-subject driven methods to a video scenario for comparison, including DreamBooth [49] and CustomDiffusion [43]. Also, we compare our method with single-subject driven video personalization method DreamVideo [11]. Moreover, we consider an alternative way for multi-subject T2V customization, that is, the combination of multi-concept T2I and I2V diffusion models. For this case, we utilize the latest multi-concept driven T2I method DisenDiff [14] for customization and I2V method SVD-XT [3] for animating the generated image. We adopt the follow settings for the comparison methods in our experiments:
| Method | Venue | CLIP-T | CLIP-I | DINO-I | T. Cons. |
| DreamBooth [49] | CVPR’23 | 0.6451 | 0.6079 | 0.3109 | 0.7084 |
| CustomDiffusion [43] | CVPR’23 | 0.6524 | 0.6206 | 0.3164 | 0.7342 |
| VideoDreamer [12] | arXiv’23 | 0.6638 | 0.6297 | 0.3479 | 0.7267 |
| DisenDiff [14] + SVD [3] | CVPR’24 | 0.6592 | 0.6179 | 0.3318 | 0.7095 |
| DreamVideo [11] | CVPR’24 | 0.6646 | 0.6128 | 0.3228 | 0.7498 |
| CustomVideo | 0.7075 | 0.6863 | 0.3983 | 0.7960 | |
| (ours vs. DreamVideo [11]) | +6.46% | +11.99% | +23.39% | +6.16% |
Evaluation Metrics
IV-B Main Results
Qualitative Results
We present the qualitative comparison results in Fig.4. From these results, we observe that our CustomVideo approach effectively ensures the co-occurrence of multiple subjects and successfully disentangles different subjects. However, both DreamBooth and CustomDiffusion fail to capture the structural color information of the car provided in our experiments. The dominant black color from the car window obscures the entire car, resulting in low fidelity. Moreover, the generated frames from VideoDreamer lack consistency, as some frames depict two cars while others only show one car. Additionally, VideoDreamer fails to capture accurate color information, such as the color of the car door. The alternative approach, DisenDiff + SVD-XT can not neither retain the identities of the given subjects nor generate videos with rich motions. Without specific design for tackling multiple subjects, DreamVideo can not distinguish the similar pair ‘cat’ and ‘dog’. In contrast, our CustomVideo method excels in handling such challenging scenes and foreground subject scenarios, effectively capturing the intricate structural details of the provided car for video generation. Similarly, when considering the case of ‘cat’ and ‘dog’, our approach also demonstrates superior capability in generating high-quality videos. More results can be found in Fig.5 and Fig.6.
Quantitative Results
We conduct quantitative experiments on our collected dataset. To ensure a thorough analysis, we generate videos using 10 individual prompts and 4 random seeds for each pair of subjects, resulting in a total of 2,720 generated videos for each method. We then evaluate the quality of the generated videos using four metrics, and the results are presented in Table IV. The table clearly demonstrates that our proposed method is capable of generating videos that are better aligned with the given subjects, outperforming the most recent video diffusion model, DreamVideo, by 11.99% and 23.39% in terms of CLIP Image Alignment and DINO Image Alignment, respectively. Meanwhile, the videos generated from our approach achieve better alignment with the desired text prompts, exceeding DreamVideo by 6.46% CLIP Textual Alignment. These improvements can be attributed to our specially designed co-occurrence and attention control mechanisms, which effectively disentangle and preserve the fidelity of the subjects. Furthermore, our CustomVideo generates videos with significantly higher temporal consistency compared to SOTA methods, as indicated in Table IV. For instance, CustomVideo surpasses DreamVideo by 6.16% in terms of Temporal Consistency. Moreover, CustomVideo outperforms DisenDiff+SVD-XT by +11.07% in CLIP Image Alignment and +12.19% in Temporal Consistency, demonstrating the effectiveness and robustness of our method in preserving the identity and modeling the motion.
Human Preference Study
To further validate our method, we conduct human evaluations on our CustomVideo with comparison to 5 SOTA methods. In this study, we collect 1500 answers from 25 independent human raters with the following questions: 1) which one is aligned to the text prompt best? 2) which one is aligned to the subject images best? 3) which one has the best overall quality? The results are shown in Fig.7. Our CustomVideo is found to be the most preferred option based on human evaluations in all three dimensions. The human annotators are all researchers with experience in artificial intelligence. The interface for human evaluation is shown in Fig.8.
IV-C Ablation Studies
We conduct ablation analysis of our method from 3 aspects: 1) the effect of each component in our method; 2) the behavior of attention control; 3) the effect of cross-attention maps from different layers; 4) the effect of hyper-parameters in attention control.
Component Analysis
We conduct a thorough analysis of each component in our CustomVideo, presenting both qualitative and quantitative results in Fig.10 and Table V. One crucial observation is the significance of ensuring co-occurrence during the fine-tuning process. We notice a significant drop in performance when multiple subjects are not concatenated into a single image (line ‘w/o concat’ in Table V). The absence of subject concatenation leads to the domination of one single subject, which is deemed unacceptable (line ‘w/o concat’ in Fig.10). Moreover, we conduct a study on fine-tuning T2V model with both single and concatenated subjects to examine if the single subject could aid in learning corresponding characteristics. Surprisingly, the results reveal that adding the single subjects to the training process proved detrimental, resulting in inconsistent generation during inference (line ‘both single and concat’ in Fig.10).
| Method | CLIP-T | CLIP-I | DINO-I | T. Cons. |
| w/o remove bg | 0.6682 | 0.6752 | 0.3636 | 0.7609 |
| w/o concat | 0.6115 | 0.6611 | 0.3386 | 0.7034 |
| both single and concat | 0.6248 | 0.6096 | 0.3054 | 0.7604 |
| w/o pos. attn. | 0.6851 | 0.6158 | 0.3215 | 0.7213 |
| w/o neg. attn. | 0.6961 | 0.6716 | 0.3613 | 0.7886 |
| CustomVideo (ours) | 0.7075 | 0.6863 | 0.3983 | 0.7960 |
Our attention control mechanism plays a crucial role in preserving the identities of subjects. The positive attention guidance in particular improves the CLIP Image Alignment metric by 11.45%, as demonstrated in Table V. The negative attention guidance also proved beneficial in promoting better image alignment. Moreover, the attention control mechanism significantly enhances the temporal consistency of the generated videos, a crucial factor in video generation. The videos generated without positive and negative attention guidance clearly showcase the positive impact of both components (lines ‘w/o pos. attn.’ and ‘w/o neg. attn.’ in Fig.10). The positive guidance specifically preserves the unique characteristics of a subject, such as distinguishing between a cat and a dog, while the negative guidance weakens the influence of other subjects on a specific subject. By incorporating both positive and negative guidance, our CustomVideo excels in generating videos with high subject fidelity and remarkable temporal consistency. Furthermore, we observe that removing the background from the given subject images significantly improves the generation of high-quality videos (line ‘w/o remove bg’ in Fig.10). By eliminating the background, the T2V model can focus solely on learning the characteristics of the given subjects.
Comparison of Cross-attention Maps
We investigate the cross-attention maps during training, and study the behavior of our attention control mechanism. The visualization results are shown in Fig.9. Without attention control, it can be observed that the learnable word token ‘new1’ for cat is aligned to the region of dog, which is unacceptable. In contrast, equipped with attention control mechanism, the learnable word token ‘new1’ and ‘new2’ are better aligned to the correct areas of cat and dog, respectively. Such results strongly demonstrate the effectiveness of our approach.
Layers of Cross-attention Maps
We investigates the effect of how we extract the cross-attention maps from the cross-attention layers. There are four levels of cross-attention layers in the U-Net, in which the sizes cross-attention maps are different. Taking the resolution of our generated low resolution video as an example, the sizes of cross-attention maps are (), (), (), and (). Here, we study the effect of different levels of cross-attention maps, the results are shown in Table VI. To tackle the size mismatch between the cross-attention maps and ground truth mask, we systematically downsample the ground truth mask to precisely match the spatial dimensions of different cross-attention maps. Specifically, when utilizing cross-attention maps from both () and (), the ground truth mask is resized to the respective size levels of and simultaneously. Attention control is then performed at both and levels to evaluate their combined impact on subject disentanglement and feature learning. We observe a significant performance drop with too large () or small () cross-attention maps. Large cross-attention maps may, on one hand, lead to the loss of subtle subject features, while on the other hand, small cross-attention maps can result in inaccurate optimization between the maps and the downsampled ground truth masks. Regarding the cross-attention maps with middle size from and , works better. Thus, in our experiments, we extract cross-attention maps from for attention control.
| CLIP-T | CLIP-I | DINO-I | T. Cons. | ||||
| ✓ | ✗ | ✗ | ✗ | 0.6111 | 0.6335 | 0.3964 | 0.7315 |
| ✗ | ✓ | ✗ | ✗ | 0.6080 | 0.6641 | 0.3055 | 0.7242 |
| ✗ | ✗ | ✓ | ✗ | 0.7075 | 0.6863 | 0.3983 | 0.7960 |
| ✗ | ✗ | ✗ | ✓ | 0.6134 | 0.6420 | 0.3215 | 0.7793 |
| ✓ | ✓ | ✗ | ✗ | 0.6234 | 0.6375 | 0.3497 | 0.7443 |
| ✗ | ✗ | ✓ | ✓ | 0.7047 | 0.6007 | 0.3642 | 0.7380 |
| ✗ | ✓ | ✓ | ✗ | 0.6730 | 0.6364 | 0.3977 | 0.7433 |
| ✓ | ✗ | ✗ | ✓ | 0.6508 | 0.6360 | 0.3055 | 0.7469 |
| ✓ | ✓ | ✓ | ✓ | 0.6432 | 0.6738 | 0.3594 | 0.6975 |
Effect of Hyper-parameters in Attention Control
We also conduct investigations into the effects of two important parameters: the weight of attention loss () and the negative value () used in the guidance mask (). The quantitative results of these investigations are presented in Table VII and Table VIII. We observe that the best performance is achieved when the weight is set to 0.2. This value results in optimal alignment between the generated videos and the given subjects, indicating the importance of appropriately balancing the attention guidance during training. For the negative value in the guidance mask, our findings reveal that even a slight negative value is sufficient to enhance the quality of T2V generation. This implies that incorporating negative attention guidance can effectively suppress the influence of other subjects on a specific subject, improving the generation quality.
IV-D Limitations
In Fig.11, we present some failure cases encountered during our experiments. Since our method relies on the ability of the base model, it would fail if the base model could not generate, such as small faces in a global view. Currently, Our approach can not tackle too many subjects, in which the spatial positions of different objects in the video conflict with each other. To address such case, we can control the positions of different subjects via conditioning on the pre-defined spatial bounding boxes during inference sampling.
| CLIP-T | CLIP-I | DINO-I | T. Cons. | |
| 1.0 | 0.6909 | 0.6496 | 0.3773 | 0.7812 |
| 0.2 | 0.7075 | 0.6863 | 0.3983 | 0.7960 |
| 0.01 | 0.6924 | 0.6710 | 0.3715 | 0.7886 |
| CLIP-T | CLIP-I | DINO-I | T. Cons. | |
| -1e-5 | 0.6786 | 0.6392 | 0.3719 | 0.7604 |
| -1e-8 | 0.7075 | 0.6863 | 0.3983 | 0.7960 |
| -1e-11 | 0.6926 | 0.6737 | 0.3613 | 0.7812 |
V Conclusions